Short Report – QA/QC Meeting Mar 2018

Date: 2018-03-27

Participants / Excused

LH, KF, RM, WE, PM, CE, MG (7) / SB, EPL

Talking Points


  1. Please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  2. SRP & Tech-SRP: please send short statement about site to LH (max. 2 sentences)
  3. Please be patient with QV changes, priority conflicts w/ ICOS.
  4. QuickViews Wishlist: what do you want for the QuickViews? Please send your wishes to LH.
  5. LH checks if possible to get own volume for temporary flux calcs if disk space is not enough. For now, LH will keep an eye on available disk space and check for separate volume if necessary.

General Issues

  1. Meteo screening tool maybe not fully compatible with flux processing tools, needs check. LH will check.


  1. [UPCOMING] Automatic fluxes will be activated again (goal is March April 2018), priority conflicts w/ ICOS.
  2. [UPCOMING] QA/QC Weekly checks in online checklists (instead of the server Excel sheet), starting w/ CH-DAV
  3. [UPCOMING] QV File Reports: checks available files on server

QuickViews Check

LH checks if QVs were succesfully generated, at least once a month, for all sites.

Problems for QVs generation

  • Windows 10 on VPC updates, shuts down the VPS and thus logs out the QV user, QVs generation stops

Generation Checklist, 26 Mar 2018

CH-AWS !not updated since Feb 2018, but current meteo data is on our server Solved.
CH-DAV !not updated since 6 Mar 2018, GL server was not reachable by DAV-MAIN PC computer. Solved, we now use a different method, in-line with all other sites, to “get” the data instead of sending.

Site Reports


SRP (LH, fluxes): So far we have nearly no gaps for fluxes in 2018, only some missed data on 3 Jan 2018. In recent days a clear diel cycle emerged for LE, while CO2 fluxes are still hovering around zero, with a tendency to emission.
Tech-SRP (IB): All instruments work fine, only the mast is not moving as it should, therefore the mast height is not adapted to the snowheight. We changed it while doing some tests on Feb. 15th. The mastheight is now at 71cm and it still does not move automatically, also moving it remotely from here is not possible.


SRP (KF): We had no instrument failure during the last month, and the laser box is renewed at the date of the meeting. Interestingly, N2O fluxes peaked the days following the 27th February, showing surprisingly high fluxes during a few days. The farmer confirmed that no management took place on the experiment and surrounding parcels.
Tech-SRP (PL): coming soon


SRP (MG, LH): Fluxes before 7 Mar 2018 look typical for winter (there is still lots of snow), with CO2 fluxes hovering around zero with a tendency to emission. The diel cycle for LE is already clearly visible, with LE increasing before the dismounting of the LI-7200. We were able to collect some more QCL fluxes before the dismounting of the QCL on 21 Mar 2018, but there were several days of data gaps before that date.
Tech-SRP (PM): LI-7200 unmounted on 7 Mar 2018 and shipped to Licor Germany for factory calibration. QC-Tildas unmounted on 21 Mar 2018 in order to ship it to Aerodyne for repairs and maintenance. Hence, GHG-Flux data incomplete.


SRP (LH): Perfect data coverage so far this year, CO2 fluxes are near-neutral. Like at other sites, the LE diel cycle is already clearly visible (also already earlier in the year).
Tech-SRP (PL): coming soon


SRP (EPL): The fluxes for CH-LAE and CH-LAS look normal for this time of the year. There was a change in data acquisition script on January 31, where the LI-7500 was removed from the raw data stream, and therefore a different EddyPro setup is needed to process the fluxes after that time. In addition, the full 2017 year has been processed for CH-LAE and CH-LAS.
Tech-SRP (TB): Working as usual


SRP (CE): The site is running generally fine, only the sonic temperature is showing spikes from time to time, which does not seem to influence the fluxes.
Tech-SRP (TB): TH3 soil temp stick is defect and replacement is ordered. R3-50 still shows strange raw data usually when raining.


You are here: Short Report - QA/QC Meeting Mar 2018


Last Updated on 30 May 2018 15:49

Leave a Reply

You must be logged in to post a comment.